# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ Abstract
There are a wide range of first passage time (FTP) problems in the physical and life sciences that can be modelled in terms of a Brownian particle binding to a reactive surface (absorption). However, prior to absorption, the particle may undergo several rounds of surface attachment (adsorption), detachment (desorption) and diffusion. Alternatively, the surface may be stochastically gated so that absorption can only occur when the gate is open. In both cases one can view each return to the surface as a renewal event. In this paper we develop a renewal theory for stochastically gated FTP problems along analogous lines to previous work on adsorption/desorption processes. We proceed by constructing a renewal equation that relates the joint probability density for particle position and the state of a gate (or multiple gates) to the probability density and FPT density for a totally absorbing (non-gated) boundary. This essentially decomposes sample paths into an alternating sequence of bulk diffusion and instantaneous adsorption/desorption events, which is terminated when adsorption coincides with an open gate. Through a variety of examples, we show how renewal theory provides a general mathematical framework for incorporating stochastic gating into FTP problems.
@@ 1 Introduction
A major topic in single-particle diffusion is the first passage time (FPT) problem for a sample path to reach a target ${\mathcal{U}}\subset\Omega$ (or a set of targets) within a bounded domain $\Omega\subset{\mathbb{R}}^{d}$ [1], see Fig. 1(a). This is often modelled in terms of Brownian motion in $\Omega\backslash{\mathcal{U}}$ supplemented by the stopping condition that the process is terminated as soon as the particle reaches the target boundary $\partial{\mathcal{U}}$. In other words, $\partial{\mathcal{U}}$ is treated as a totally absorbing surface. (Note that the exterior boundary $\partial\Omega$ could also be an absorbing target, and there could be multiple interior targets. Here $\partial\Omega$ is assumed to be totally reflecting.) Absorption may represent a particle binding to the surface and triggering a downstream event such as a chemical reaction or the transfer of resources between the particle and target. Examples include an extracellular signalling molecule (ligand) binding to a receptor in the surface membrane of a cell [2], and a foraging animal finding a target containing food or shelter [3, 4, 5]. However, the assumption of instantaneous (total) absorption neglects the possibility that when the particle reaches the target surface, additional constraints must be satisfied before the stochastic process is terminated.
@@
One consequence of a sticky boundary $\partial{\mathcal{U}}$ is that a particle typically undergoes several rounds of bulk diffusion followed by first-return to $\partial{\mathcal{U}}$ prior to a final absorption event. Each first-return to $\partial{\mathcal{U}}$ can be viewed as a renewal of the stochastic process. This has motivated the construction of renewal equations that relate the probability density and FPT density for absorption to the corresponding quantities for adsorption, which are usually simpler to compute [44, 43, 45, 46]. The renewal equations take the form of implicit integral equations that can be solved explicitly using Laplace transforms and the convolution theorem. However, a crucial step in the analysis is ensuring that diffusion restarts in a state that avoids immediate re-adsorption. This can be achieved by taking the sticky surface to be partially adsorbing [44, 43, 45] or by resetting the particle to its initial position away from the wall following desorption [46]. This form of desorption-induced stochastic resetting is distinct from spontaneous resetting in the bulk domain [47]. The renewal approach provides a general framework for incorporating more complicated models of adsorption and desorption. First, one can consider a non-Markovian model of desorption by taking the waiting time density for the duration of an adsorbed state to be non-exponential. Second, one can take the target surface to be partially rather than totally adsorbing and apply encounter-based methods [44, 45]. Third, one can include both spontaneous stochastic resetting in the bulk and desorption-induced resetting [46].
In this paper we develop an analogous renewal theory for stochastically gated diffusion processes. Suppose that the target surface $\partial U$ in Fig. 1(a) switches between an absorbing boundary (open gate) and a reflecting boundary (closed gate) according to a two-state Markov chain. For the moment we do not distinguish between a particle such as an enzyme switching between a reactive and a non-reactive state, and the boundary itself being physically gated. In our renewal formulation we treat each particle encounter with $\partial{\mathcal{U}}$ when the gate is closed as an instantaneous adsorption/desorption event. The corresponding renewal equation relates the joint probability density for particle position and the state of the gate (open or closed) to the probability density and FPT density for a totally absorbing (non-gated) surface. That is, sample paths are effectively decomposed into an alternating sequence of bulk diffusion and instantaneous adsorption/desorption events, which is terminated when adsorption coincides with an open gate. As in the case of sticky targets, we need to restart the diffusion process in a state that avoids immediate re-adsorption. Therefore, we assume that following adsorption at a point $\mathbf{y}\in\partial{\mathcal{U}}$, the particle immediately restarts at a position $\mathbf{a}_{\epsilon}(\mathbf{y})$ such that $\lim_{\epsilon\rightarrow 0}\mathbf{a}_{\epsilon}(\mathbf{y})=\mathbf{y}$. The integral renewal equation is solved explicitly in Laplace space for $\epsilon>0$ and the solution of the original problem is then obtained by taking the limit $\epsilon\rightarrow 0$. Through a variety of examples, we show how the renewal formulation provides a general mathematical framework for incorporating stochastic gating into FPT problems.
@@
The structure of the paper is as follows. In section 2 we introduce the renewal approach by considering a particle diffusing on the half-line $0,\infty)$ with a stochastically gated boundary at $x=0$. Assuming that the particle instantaneously resets to a position $\epsilon>0$ whenever it is reflected off the closed gate, we solve the associated renewal equation in Laplace space. We then show how taking the limit $\epsilon\rightarrow 0$ recovers the solution obtained by directly solving the forward Kolmogorov equation for the switching system. In section 3, we apply the renewal method to a pair of 1D FPT problems. First, we calculate the MFPT for absorption of a Brownian particle with stochastic resetting on the half-line and a stochastically gated boundary. Second, we consider the FPT problem for single-particle diffusion in the finite interval with a stochastically gated boundary at each end. If the particle switches between a reactive and non-reactive state then both ends are either totally absorbing or totally reflecting (totally correlated boundary conditions). On the other hand, if the boundaries themselves are stochastically gated, then they independently switch between open and closed states (uncorrelated boundary conditions). In section 4, we consider a natural extension of the 1D problems analysed in section 3, namely, diffusion in a domain $\Omega\subset{\mathbb{R}}^{d}$, $d>1$, with a stochastically gated target surface $\partial\Omega\subset{\mathbb{R}}^{d-1}$. The corresponding renewal equation now involves a spatial integral with respect to $\mathbf{y}\in\partial\Omega$, which complicates the analysis. Nevertheless, we show how the Laplace transformed renewal equation can be solved by extending spectral methods previously developed in Refs. [8, 11, 44, 45]. In particular, taking the limit $\epsilon\rightarrow 0$ naturally leads to the appearance of a Dirichlet-to-Neumann (D-to-N) operator on the target surface $\partial\Omega$. We illustrate the theory using the example of diffusion in a sphere.
@@
We end this introduction by noting that a complementary renewal formulation of stochastically gated diffusion processes (and discrete processes on networks) has recently been developed by Scher et al. [21, 22, 23, 24]. These authors focus on a stochastically gated trap $\mathcal{U}$} in 1D rather than a point-like boundary $\partial{\mathcal{U}}$ – a particle can freely enter and exit the trap and is absorbed within $\mathcal{U}$} as soon as the gate opens. Rather than constructing integral renewal equations for the full propagator, they derive recursive equations relating the FPT density for absorption to the corresponding FPT density for adsorption in the absence of gating by conditioning on whether the particle is located within or outside the trap when the gate reopens. This allows the authors to obtain exact expressions for the Laplace transform of the FPT density. In the discussion, we briefly indicate how to extend our renewal approach to a stochastically gated trap $[a,b]$ in the half-line. The resulting renewal equation for the propagator involves a spatial integral with respect to points $y\in[a,b]$ that complicates the analysis. (Avoiding such spatial integrals is one of the motivations for the renewal approach of Scher et al. [21, 22, 23, 24].) Nevertheless, the solution can be expressed in terms of a Neumann series expansion along analogous lines to adsorption/desorption process within traps [45]. Higher-dimensional versions can also be constructed.
@@
2 Brownian particle on the half-line with a
stochastically gated boundary
@@ 2.1 Forward Kolmogorov equation
${\mathcal{F}}_{k_{0}}(x_{0},t)\equiv-\frac{d{\mathfrak{Q}}_{k_{0}}(x_{0},t)}{dt}=D\frac{\partial\rho_{1|k_{0}}(0,t|x_{0})}{\partial x}.$ (2.6)
@@
$D\frac{\partial^{2}\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})}{\partial x^{2}}-(\alpha+s)\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})+\beta\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})=-\delta(x-x_{0})\delta_{1,k_{0}},$ (2.7a)
$D\frac{\partial^{2}\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})}{\partial x^{2}}-(\beta+s)\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})+\alpha\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})=-\delta(x-x_{0})\delta_{-1,k_{0}}$ (2.7b)
@@
$D\frac{\partial\widetilde{\rho}_{-1|k_{0}}(0,s|x_{0})}{\partial x}=0,\quad\widetilde{\rho}_{1|k_{0}}(0,s|x_{0})=0.$ (2.7c)
Adding equations (2.7a,b) and setting $q(x,s)=\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})+\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})$, we have
$D\frac{\partial^{2}q(x,s)}{\partial x^{2}}-sq(x,s)=-\delta(x-x_{0}),\quad x>0.$ (2.8)
@@
$q(x,s)$ $=\widetilde{G}(x,s|x_{0})+{\mathrm{e}}^{-\sqrt{s/D}x}A(s),$ (2.9)
@@
$\widetilde{G}(x,s|x_{0})=\frac{1}{2\sqrt{sD}}\bigg({\mathrm{e}}^{-\sqrt{s/D}|x-x_{0}|}-{\mathrm{e}}^{-\sqrt{s/D}(x+x_{0})}\bigg)$ (2.10)
@@
$\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})=\sigma_{1}q(x,s)+j_{k_{0}}(x,s),\quad\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})=\sigma_{-1}q(x,s)-j_{k_{0}}(x,s),$
with $j_{k_{0}}(0,s)=B(s)$. Substituting into equation (2.7a) implies that
$D\frac{\partial^{2}j_{k_{0}}(x,s)}{\partial x^{2}}-(\alpha+\beta+s)j_{k_{0}}(x,s)=-\delta(x-x_{0})\bigg[\delta_{1,k_{0}}-\sigma_{1}\bigg],\quad x>0.$ (2.12)
@@
$j_{k_{0}}(x,s)=k_{0}\sigma_{-k_{0}}\widetilde{G}(x,s+\alpha+\beta|x_{0})-{\mathrm{e}}^{-\sqrt{(s+\alpha+\beta)/D}x}B(s),\quad x>0.$ (2.13)
We have two unknown coefficients $A(s),B(s)$, which also depend on $(x_{0},k_{0})$, and two linearly independent boundary conditions given by (2.7c). Expressing the latter in terms of $q$ and $j_{k_{0}}$, we have
$\left.\sigma_{-1}\frac{\partial q(x,s)}{\partial x}\right|_{x=0}=\left.\frac{\partial j_{k_{0}}(x,s)}{\partial x}\right|_{x=0},\quad\sigma_{1}q(0,s)+j_{k_{0}}(0,s)=0.$ (2.14)
Combing with the general solutions (2.9) and (2.13) gives $B(s)=\sigma_{1}A(s)$ with
$A(s)=\frac{\sigma_{-1}{\mathrm{e}}^{-\sqrt{s/D}x_{0}}-k_{0}\sigma_{-k_{0}}{\mathrm{e}}^{-\sqrt{(s+\alpha+\beta)/D}x_{0}}}{\sigma_{-1}\sqrt{sD}+\sigma_{1}\sqrt{(s+\alpha+\beta)D}}.$ (2.15)
In summary, given these coefficients, the solution of equations (2.7a)– (2.7c) is
@@
$\hskip 56.9055pt+k\bigg[k_{0}\sigma_{-k_{0}}\widetilde{G}(x,s+\alpha+\beta|x_{0})-\sigma_{1}{\mathrm{e}}^{-\sqrt{[s+\alpha+\beta]/D}x}A(s)\bigg].$ (2.16)

### v2

@@ Abstract
There are a wide range of first passage time (FPT) problems in the physical and life sciences that can be modelled in terms of a Brownian particle binding to a reactive surface (target) and initiating a downstream event (absorption). However, prior to absorption, the particle may undergo several rounds of surface attachment (adsorption), detachment (desorption) and diffusion. Alternatively, the surface may be stochastically gated so that absorption can only occur when the gate is open. In both cases one can view each return to the surface as a renewal event. In this paper we develop a renewal theory for stochastically gated FPT problems along analogous lines to previous work on adsorption/desorption processes. We proceed by constructing a renewal equation that relates the joint probability density for particle position and the state of a gate (or multiple gates) to the probability density and FPT density for a totally absorbing (non-gated) boundary. This essentially decomposes sample paths into an alternating sequence of bulk diffusion and instantaneous adsorption/desorption events, which is terminated when adsorption coincides with an open gate. In order to ensure that diffusion restarts in a state that avoids immediate re-adsorption. we assume that whenever the particle reaches a closed boundary it is instantaneously shifted a distance $\epsilon$ from the boundary (boundary-induced resetting). We explicitly solve the renewal equation for $\epsilon>0$ and show how the solution to the original gated FPT problem is recovered in the limit $\epsilon\rightarrow 0$. Through a series of examples, we who how renewal theory proves a general mathematical framework for modelling stochastically gated targets.
@@ 1 Introduction
A major topic in single-particle diffusion is the first passage time (FPT) problem for a sample path to reach a target ${\mathcal{U}}\subset\Omega$ (or a set of targets) within a bounded domain $\Omega\subset{\mathbb{R}}^{d}$ [1], see Fig. 1(a). This is often modelled in terms of Brownian motion (BM) in $\Omega\backslash{\mathcal{U}}$ supplemented by the stopping condition that the process is terminated as soon as the particle reaches the target boundary $\partial{\mathcal{U}}$. In other words, $\partial{\mathcal{U}}$ is treated as a totally absorbing surface. (Note that the exterior boundary $\partial\Omega$ could also be an absorbing target, and there could be multiple interior targets. Here $\partial\Omega$ is assumed to be totally reflecting.) Absorption may represent a particle binding to the surface and triggering a downstream event such as a chemical reaction or the transfer of resources between the particle and target. Examples include an extracellular signalling molecule (ligand) binding to a receptor in the surface membrane of a cell [2], and a foraging animal finding a target containing food or shelter [3, 4, 5]. However, the assumption of instantaneous (total) absorption neglects the possibility that when the particle reaches the target surface, additional constraints must be satisfied before the stochastic process is terminated.
@@
One consequence of a sticky boundary $\partial{\mathcal{U}}$ is that a Brownian particle typically undergoes several rounds of bulk diffusion followed by first-return to $\partial{\mathcal{U}}$ prior to a final absorption event. Each first-return to $\partial{\mathcal{U}}$ can be viewed as a renewal of the stochastic process. This has motivated the construction of renewal equations that relate the probability density and FPT density for absorption to the corresponding quantities for adsorption, which are usually simpler to compute [44, 43, 45, 46]. The renewal equations take the form of implicit integral equations that can be solved explicitly using Laplace transforms and the convolution theorem. However, a crucial step in the analysis is ensuring that BM restarts in a state that avoids immediate re-adsorption. This can be achieved by taking the sticky surface to be partially adsorbing [44, 43, 45] or by resetting the particle to its initial position away from the wall following desorption [46]. This form of desorption-induced stochastic resetting is distinct from spontaneous resetting in the bulk domain [47]. The renewal approach provides a general framework for incorporating more complicated models of adsorption and desorption. First, one can consider a non-Markovian model of desorption by taking the waiting time density for the duration of an adsorbed state to be non-exponential. Second, one can take the target surface to be partially rather than totally adsorbing and apply encounter-based methods [44, 45]. Third, one can include both spontaneous stochastic resetting in the bulk and desorption-induced resetting [46].
@@
In this paper we develop an analogous renewal theory for stochastically gated diffusion processes. We begin by considering BM on the half-line $0,\infty)$ with a stochastically gated boundary at $x=0$ (section 2). We first solve the forward Kolmogorov equation for the probability densities $\rho_{k}(x,t)={\mathbb{E}}[\rho(x,t)\delta_{k,\sigma(t)}]$, where $\rho(x,t)$ is the marginal probability density for particle position anmd expectations are taken with respect to realisations of the gating process. We then construct a first renewal equation that relates $\rho_{k}(x,t)$ to the corresponding probability density $p(x,t)$ for BM on the half-line with a totally absorbing boundary condition at $x=0$. In the renewal formulation, we treat each particle encounter with the boundary $x=0$ when the gate is closed as an instantaneous adsorption/desorption event. That is, sample paths are effectively decomposed into an alternating sequence of bulk diffusion and instantaneous adsorption/desorption events, which is terminated when adsorption coincides with an open gate. As in the case of sticky targets, we need to restart the diffusion process in a state that avoids immediate re-adsorption. Therefore, we assume that following adsorption at $x=0$ when the gate is closed, the particle immediately restarts at a distance $\epsilon$ from the origin. The latter can be interpreted as a form of so-called boundary-induced resetting [46]. We then obtain an explicit solution of the renewal equation for $\epsilon>0$ using Laplace transforms and the convolution theorem. We then show how the resulting solution in the limit $\epsilon\rightarrow 0$ is equivalent to the solution of the Kolmogorov equation. However, the renewal equation provides a more general mathematical framework for modelling a stochastically gated boundary by explicitly separating the FPT problem for reaching the detecting the boundary (absorption) and the subsequent rule for restarting diffusion. In section 3, we apply the renewal method to a pair of 1D FPT problems. First, we calculate the MFPT for absorption of a Brownian particle with stochastic resetting on the half-line and a stochastically gated boundary. Second, we consider the FPT problem for single-particle diffusion in the finite interval with a stochastically gated boundary at each end. If the particle switches between a reactive and non-reactive state then both ends are either totally absorbing or totally reflecting (totally correlated boundary conditions). On the other hand, if the boundaries themselves are stochastically gated, then they independently switch between open and closed states (uncorrelated boundary conditions). In section 4, we consider a natural extension of the 1D problems analysed in section 3, namely, diffusion in a domain $\Omega\subset{\mathbb{R}}^{d}$, $d>1$, with a stochastically gated target surface $\partial\Omega\subset{\mathbb{R}}^{d-1}$. The corresponding renewal equation now involves a spatial integral with respect to $\mathbf{y}\in\partial\Omega$, which complicates the analysis. Nevertheless, we show how the Laplace transformed renewal equation can be solved by extending spectral methods previously developed in Refs. [8, 11, 44, 45]. In particular, taking the limit $\epsilon\rightarrow 0$ naturally leads to the appearance of a Dirichlet-to-Neumann (D-to-N) operator on the target surface $\partial\Omega$. We illustrate the theory using the example of diffusion in a sphere.
@@
We end this introduction by noting that a complementary renewal formulation of stochastically gated processes has recently been developed by Scher et al. [21, 22, 23, 24]. In particular, in Ref. [24] these authors focus on a stochastically gated trap $\mathcal{U}$} in 1D rather than a point-like boundary $\partial{\mathcal{U}}$ – a diffusing particle can freely enter and exit the trap and is absorbed within $\mathcal{U}$} as soon as the gate opens. Instead of constructing integral renewal equations for the full propagator, they derive recursive equations relating the FPT for absorption to the corresponding FPT density for adsorption in the absence of gating by conditioning on whether the particle is located within or outside the trap when the gate reopens. This allows the authors to obtain exact expressions for the Laplace transform of the FPT density for a 1D trap. In the discussion, we briefly indicate some of the difficulties in extending our renewal formulation to the trapping problem.
@@
2 Brownian particle on the half-line with a stochastically gated boundary
@@ 2.1 Forward Kolmogorov equation
${\mathcal{F}}_{k_{0}}(x_{0},t)\equiv-\frac{d{\mathfrak{Q}}_{k_{0}}(x_{0},t)}{dt}=D\frac{\partial\rho_{1|k_{0}}(0,t|x_{0})}{\partial x}.$ (2.4f)
@@
$D\frac{\partial^{2}\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})}{\partial x^{2}}-(\alpha+s)\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})+\beta\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})=-\delta(x-x_{0})\delta_{1,k_{0}},$ (2.4ga)
$D\frac{\partial^{2}\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})}{\partial x^{2}}-(\beta+s)\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})+\alpha\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})$
$\hskip 142.26378pt=-\delta(x-x_{0})\delta_{-1,k_{0}}$ (2.4gb)
@@
$D\frac{\partial\widetilde{\rho}_{-1|k_{0}}(0,s|x_{0})}{\partial x}=0,\quad\widetilde{\rho}_{1|k_{0}}(0,s|x_{0})=0.$ (2.4gc)
Adding equations (2.4ga) and (2.4g), and setting
$q(x,s)=\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})+\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0}),$ (2.4gh)
we have
$D\frac{\partial^{2}q(x,s)}{\partial x^{2}}-sq(x,s)=-\delta(x-x_{0}),\quad x>0.$ (2.4gi)
@@
$q(x,s)$ $=\widetilde{G}(x,s|x_{0})+{\mathrm{e}}^{-\sqrt{s/D}x}A(s),$ (2.4gj)
@@
$\widetilde{G}(x,s|x_{0})=\frac{1}{2\sqrt{sD}}\bigg({\mathrm{e}}^{-\sqrt{s/D}|x-x_{0}|}-{\mathrm{e}}^{-\sqrt{s/D}(x+x_{0})}\bigg)$ (2.4gk)
@@
$\widetilde{\rho}_{1|k_{0}}(x,s|x_{0})=\sigma_{1}q(x,s)+j_{k_{0}}(x,s),\quad\widetilde{\rho}_{-1|k_{0}}(x,s|x_{0})=\sigma_{-1}q(x,s)-j_{k_{0}}(x,s),$ (2.4gl)
with $j_{k_{0}}(0,s)=B(s)$. Substituting into equation (2.4ga) implies that
$D\frac{\partial^{2}j_{k_{0}}(x,s)}{\partial x^{2}}-(\alpha+\beta+s)j_{k_{0}}(x,s)=-\delta(x-x_{0})\bigg[\delta_{1,k_{0}}-\sigma_{1}\bigg],\quad x>0.$ (2.4gm)
@@
$j_{k_{0}}(x,s)=k_{0}\sigma_{-k_{0}}\widetilde{G}(x,s+\alpha+\beta|x_{0})-{\mathrm{e}}^{-\sqrt{(s+\alpha+\beta)/D}x}B(s),\quad x>0.$ (2.4gn)
We have two unknown coefficients $A(s),B(s)$, which also depend on $(x_{0},k_{0})$, and two linearly independent boundary conditions given by (2.4gc). Expressing the latter in terms of $q$ and $j_{k_{0}}$, we have
$\left.\sigma_{-1}\frac{\partial q(x,s)}{\partial x}\right|_{x=0}=\left.\frac{\partial j_{k_{0}}(x,s)}{\partial x}\right|_{x=0},\quad\sigma_{1}q(0,s)+j_{k_{0}}(0,s)=0.$ (2.4go)
Combining with the general solutions (2.4gj) and (2.4gn) gives $B(s)=\sigma_{1}A(s)$ with
$A(s)=\frac{\sigma_{-1}{\mathrm{e}}^{-\sqrt{s/D}x_{0}}-k_{0}\sigma_{-k_{0}}{\mathrm{e}}^{-\sqrt{(s+\alpha+\beta)/D}x_{0}}}{\sigma_{-1}\sqrt{sD}+\sigma_{1}\sqrt{(s+\alpha+\beta)D}}.$ (2.4gp)
In summary, given these coefficients, the solution of equations (2.4ga)– (2.4gc) is
@@
$\hskip 56.9055pt+k\bigg[k_{0}\sigma_{-k_{0}}\widetilde{G}(x,s+\alpha+\beta|x_{0})-\sigma_{1}{\mathrm{e}}^{-\sqrt{[s+\alpha+\beta]/D}x}A(s)\bigg].$ (2.4gq)